# ALL ORDINALS ARE COP-ROBBER ORDINALS

### v2

@@ 1 Introduction
Let $G$ be a fixed graph. Two players, the cop and the robber, each pick a starting vertex, with the cop picking first. Then they move alternately, with the cop moving first: at each turn the player moves to an adjacent vertex or does not move. The game is won by the cop if he lands on the robber. We say that $G$ is cop-win if the cop has a winning strategy.
A central notion for this problem is that of a ‘dominated’ vertex. We say that a vertex $y$ dominates a vertex $x$ if the set of $x$ and all neighbors of $x$, denoted by $N[x]$, is contained in the set of $y$ and all neighbors of $y$. The finite cop-win graphs were characterised by Nowakowski and Winkler [5]. They showed that a finite graph $G$ is cop-win if and only if it is constructible, meaning that it can be built up from the one-point graph by repeatedly adding dominated vertices. There are multiple variations of this game for finite graphs, such as allowing more than one cop to play. We direct the interested reader to the book of Bonato and Nowakowski [1] for general background and a wealth of other results in the finite case.
We now turn to infinite graphs, where the game of cops and robbers has the exact same rules as before, but the above notions behave very differently. For example, an infinite ray is constructible, but it is trivially not a cop-win graph. Conversely, there exist cop-win graphs that are not constructible – for this, and some other phenomena, see Ivan, Leader and Walters [4]. We mention that currently, unlike the finite graphs setup, there is no characterization for infinite cop-win graphs.
From now on, all the graphs will be assumed to be cop-win. For a finite graph, we can define the capture time to be the length of the game, assuming both players play optimally. That means that the cop chooses his starting position as favorably as possible, and so does the robber. After that, every cop’s next move will be to minimise the time he needs to capture the robber, and every robber’s next move will be to escape for as long as possible. For example, if $G=P_{2n+1}$, the capture time is $n$. How can we generalise this notion for infinite graphs?
@@
Let $G$ be a graph of cardinality $\aleph_{\beta}$, where $\beta\geq 0$ is an ordinal. For any vertex $v\in V(G)$ we denote by $N[v]$ its closed neighborhood (its neighbors, together with $v$). Set $\omega(G)=\omega_{\beta+1}$, and define the relations $\{\leq\alpha\}_{\alpha<\omega(G)}$ on $V(G)$ as follows. If $u=v$, then $u\leq_{0}v$. Also, $u\leq_{\alpha}v$ if for all $x\in N[u]$, there exists $y\in N[v]$ such that $x\leq_{\gamma}y$ for some $\gamma<\alpha$.
@@ 2 Maximum capture time as a limit ordinal
Let the cop be at some vertex $v_{0}$. From there, in his first move, he can go to a vertex on the diagonal, which we call $v_{1}$. Since any two vertices on the diagonal are connected, the robber at his original position $u_{0}$ has to move to a vertex $u_{1}=(\alpha_{1},\beta_{1})$ such that $\alpha_{1}\neq\beta_{1}$. Without loss of generality, we may assume that $\alpha_{1}>\beta_{1}$. Next, the cop moves to a vertex $v_{2}=(\xi_{2},0)$ with $\xi_{2}>\alpha_{1}$. In order to evade the cop, the robber must move to a vertex $u_{2}=(\alpha_{2},\beta_{2})$ outside of $N[v_{2}]$. By construction, if $u_{2}\in N[u_{1}]\setminus N[v_{2}]$, then $\alpha_{2}\geq\xi_{2}$ and $\beta_{2}<\beta_{1}$ (which in particular means that $\alpha_{2}>\beta_{2}$). Continuing with this strategy, with the cop always staying on the $x$-axis, he can force the robber to always move down and to the right, decreasing his $y$-coordinate, and thus creating a strictly decreasing sequence of ordinals. Since there does not exist an infinite strictly decreasing sequence of ordinals, there exists a finite moment when the robber must reach the $x$-axis, where he will be caught in the next move. ∎

### v3

@@ 1 Introduction
Let $G$ be a fixed simple graph without multiple edges. Two players, the cop and the robber, each pick a starting vertex, with the cop picking first. Then they move alternately, with the cop moving first: at each turn the player moves to an adjacent vertex or does not move. The game is won by the cop if he lands on the robber. We say that $G$ is cop-win if the cop has a winning strategy.
A central notion for this problem is that of a ‘dominated’ vertex. For a vertex $v\in V(G)$ we denote by $N[v]$ its closed neighborhood: its neighbors, together with $v$. We say that a vertex $y$ dominates a vertex $x$ if $N[x]\subseteq N[y]$. The finite cop-win graphs were characterised by Nowakowski and Winkler [5]. They showed that a finite graph $G$ is cop-win if and only if it is constructible, meaning that it can be built up from the one-point graph by repeatedly adding dominated vertices. There are multiple variations of this game for finite graphs, such as allowing more than one cop to play. We direct the interested reader to the book of Bonato and Nowakowski [1] for general background and a wealth of other results in the finite case.
We now turn to infinite graphs, where the game of cops and robbers has the exact same rules as before, but the above notions behave very differently. For example, a ray – infinite path with one end – is constructible, but it is trivially not a cop-win graph. Conversely, there exist cop-win graphs that are not constructible – for this, and some other phenomena, see Ivan, Leader and Walters [4]. We mention that currently, unlike the finite graphs setup, there is no structural characterization for infinite cop-win graphs.
From now on, all the graphs will be assumed to be cop-win. For a finite graph, we can define the capture time to be the length of the game, assuming both players play optimally. That means that the cop chooses her starting position as favorably as possible, and so does the robber. After that, every cop’s next move will be to minimise the time she needs to capture the robber, and every robber’s next move will be to escape for as long as possible. For example, if $G=P_{2n+1}$, the capture time is $n$. How can we generalise this notion for infinite graphs?
@@
Let $G$ be a graph of cardinality $\aleph_{\beta}$, where $\beta\geq 0$ is an ordinal. Set $\omega(G)=\omega_{\beta+1}$, and define the relations $\{\leq\alpha\}_{\alpha<\omega(G)}$ on $V(G)$ as follows. If $u=v$, then $u\leq_{0}v$. Also, $u\leq_{\alpha}v$ if for all $x\in N[u]$, there exists $y\in N[v]$ such that $x\leq_{\gamma}y$ for some $\gamma<\alpha$.
@@ 2 Maximum capture time as a limit ordinal
Let the cop be at some vertex $v_{0}$. From there, in her first move, she can go to a vertex on the diagonal, which we call $v_{1}$. Since any two vertices on the diagonal are connected, the robber at his original position $u_{0}$ has to move to a vertex $u_{1}=(\alpha_{1},\beta_{1})$ such that $\alpha_{1}\neq\beta_{1}$. Without loss of generality, we may assume that $\alpha_{1}>\beta_{1}$. Next, the cop moves to a vertex $v_{2}=(\xi_{2},0)$ with $\xi_{2}>\alpha_{1}$. In order to evade the cop, the robber must move to a vertex $u_{2}=(\alpha_{2},\beta_{2})$ outside of $N[v_{2}]$. By construction, if $u_{2}\in N[u_{1}]\setminus N[v_{2}]$, then $\alpha_{2}\geq\xi_{2}$ and $\beta_{2}<\beta_{1}$ (which in particular means that $\alpha_{2}>\beta_{2}$). Continuing with this strategy, with the cop always staying on the $x$-axis, she can force the robber to always move down and to the right, decreasing his $y$-coordinate, and thus creating a strictly decreasing sequence of ordinals. Since there does not exist an infinite strictly decreasing sequence of ordinals, there exists a finite moment when the robber must reach the $x$-axis, where he will be caught in the next move. ∎